The function of y= (3x^2+x-2) / (e^x) has a horizontal tangent line when x=?
Answer #1

Find the derivative as

2 er

\mathrm{Apply\:the\:Quotient\:Rule}:\quad \left(\frac{f}{g}\right)^{'}=\frac{f\:'\cdot g-g'\cdot f}{g^2}

=\frac{\frac{d}{dx}\left(3x^2+x-2\right)e^x-\frac{d}{dx}\left(e^x\right)\left(3x^2+x-2\right)}{\left(e^x\right)^2}

(er)

=\frac{-3x^2+5x+3}{e^x}

Now, horizontal tangent line means, slope is 0, hence set numerator to 0

-3x^2+5x+3=0

\mathrm{For\:}\quad a=-3,\:b=5,\:c=3:\quad x_{1,\:2}=\frac{-5\pm \sqrt{5^2-4\left(-3\right)3}}{2\left(-3\right)}

x = \frac { 5 } { 6 } - \frac { \sqrt { 61 } } { 6 },\frac { 5 } { 6 } + \frac { \sqrt { 61 } } { 6 }